### v2

@@ Abstract
Pneumonia, a widespread respiratory infection, remains a leading cause of morbidity and mortality globally, especially in vulnerable populations. Chest X-rays are a primary tool in detecting pneumonia; however, differences in imaging conditions and subtle visual indicators can make consistent interpretation challenging. Automated tools can complement traditional methods by providing additional insights to enhance diagnostic reliability and support clinical decision-making. In this study, we propose a novel approach that leverages deep learning with integrated self-attention mechanisms to enhance pneumonia detection in chest X-rays. Our method begins with lung segmentation using a TransUNet model that integrates our specialized transformer module, which has fewer parameters compared to common transformers while maintaining performance. This model is trained on the ”Chest Xray Masks and Labels” dataset and then applied to the ”Kermany” and ”Cohen” datasets to generate lung regions, enhancing subsequent classification tasks. By isolating the lung areas, the segmentation step provides focused input for the classification stage, enabling the model to concentrate on relevant regions. For classification, we utilize pre-trained ResNet models (ResNet-50 and ResNet-101) specifically to extract multi-scale feature maps, a core aspect of our approach. These feature maps are then processed through our modified transformer module to enhance pneumonia detection. By employing our specialized transformer, we achieve superior results with significantly fewer parameters compared to common transformer models. Our approach achieves high accuracy rates of 92.79% on the ”Kermany” dataset and 95.11% on the ”Cohen” dataset, ensuring robust and efficient performance suitable for resource-constrained environments. https://github.com/amirrezafateh/Multi-Scale-Transformer-Pneumonia
@@ 1 Introduction
A serious respiratory condition known as pneumonia results in inflammation of one or both lungs, which can cause fever, coughing, and breathing difficulties. Since it causes about 15% of mortality in children under five, it is especially risky for young children unicef_pneumonia . This illness is widespread in developing and impoverished nations, where access to healthcare is restricted, and situations like pollution, overcrowding, and subpar living conditions worsen the situation who_pneumonia_in_children .
In order to effectively treat pneumonia and improve patient outcomes, early diagnosis is essential. Still, identifying pneumonia can be challenging. It can be easily confused with other lung diseases, and making consistent and accurate diagnosis difficult health_poverty_action_poverty_and_health . Conventional imaging methods are frequently employed, including computed tomography (CT), magnetic resonance imaging (MRI), and chest radiographs. But, chest X-rays are favored for being non-invasive and cost-effective radiologyinfo_chest_xray .
The variability in interpreting chest X-ray images underscores the need for consistent, automated diagnostic tools. The development of deep learning technology offers a promising approach to addressing the challenges in pneumonia diagnosis gayathri2022computer ; lin2023enhancing ; ortiz2022automatic . As a branch of artificial intelligence (AI), deep learning entails training neural networks on extensive datasets to identify patterns and generate predictions. In medical imaging, deep learning algorithms can scrutinize chest X-rays with remarkable precision, often exceeding human radiologists in terms of consistency and speed. Convolutional Neural Networks (CNNs), a specialized deep learning model designed for image analysis, have been developed to identify pneumonia from chest X-rays askari2024enhancing . These models are trained to recognize patterns and subtle features associated with pneumonia, contributing to consistent and reliable diagnostic performance that complements the expertise of radiologists krizhevsky2012imagenet ; parhami2023comparison . However, traditional CNN-based models still cannot achieve impressive accuracy. Recent advancements include the use of ensemble learning techniques, which combine predictions from multiple models to improve overall accuracy, and the integration of attention mechanisms, which help models focus on the most relevant parts of the image fateh2024advancing . Transformers have demonstrated significant potential in visual tasks. They excel at modeling long-range dependencies and identifying intricate patterns in medical images fateh2024msdnet ; mabrouk2022pneumonia .
Lung segmentation is frequently used in chest X-ray images to improve the accuracy of pneumonia detection. This technique isolates the lung regions, allowing segmentation algorithms to minimize noise and concentrate the analysis on the pertinent areas. The U-Net architecture, known for its encoder-decoder structure and skip connections, has been particularly effective in medical image segmentation, including lung segmentation in chest X-rays. Accurate segmentation allows subsequent classification models to better identify pneumonia, leading to improved diagnostic performance rajpurkar2017chexnet . Furthermore, advancements in segmentation techniques, such as integrating attention mechanisms and transformers, have improved segmentation accuracy. These improvements enable models to concentrate on the most critical parts of the lung, capturing intricate details and ensuring that even the subtle signs of pneumonia are detected zhang2023cams ; fateh2023enhancing . Effective lung segmentation thus provides a robust foundation for classification algorithms, enabling them to achieve higher accuracy and reliability in diagnosing pneumonia.
A significant challenge in training these deep learning models is the requirement for large, high-quality labeled datasets, which are often hard to acquire in the medical field. Transfer learning tackles this problem by utilizing models pre-trained on extensive datasets like ImageNet and then fine-tuning them for particular medical tasks using smaller datasets. This approach has significantly improved the accuracy of medical image classifications chen2017rethinking . Even with transfer learning, models trained in one domain often fail to perform well in significantly different domains, such as medical applications. This highlights the need for more advanced and higher-level features, which is why transformers are increasingly used rezvani2024fusionlungnet . Despite the benefits of using transformers and their achievements in segmentation and classification tasks in CXR images, they bring two new challenges: models are more complex and have many parameters. These complex models used more hardware resources and took longer to train.
@@
We propose an innovation approach that leverages deep learning with integrated self-attention mechanisms to overcome the challenges of traditional transformers while maintaining lower model complexity. Our method begins with lung segmentation using a TransUNet model, which integrates transformer-based attention mechanisms into the U-Net architecture. The TransUNet model is trained on the ”Chest Xray Masks and Labels” dataset candemir2013lung ; jaeger2013automatic to accurately segment lung regions in the images. Once trained, this pre-trained model is used with frozen weights to predict lung masks for our target datasets, ”Kermany” kermany2018labeled and ”Cohen” cohen2020covid . This segmentation step isolates the lung regions, enhancing the subsequent classification task.
@@
  • Introduction of a novel TransUNet architecture for the segmentation task.
@@
  • Achieving high accuracy rates of 92.79% on the ”Kermany” dataset and 95.11% on the ”Cohen” dataset.
@@ 2 Related Work
In recent years, the focus of research on diagnosing and categorizing lung diseases, including pneumonia, through medical imaging has intensified, driven by advances in machine learning and deep learning technologies. Precisely segmenting lung areas in chest X-ray (CXR) images is essential for reliable disease identification and thorough analysis. This section examines deep learning techniques for segmenting and diagnosing lung diseases in chest X-ray (CXR) images. For the segmentation task, we focus on the U-Net architecture and its variations, including attention mechanisms and transformer blocks, which have significantly advanced lung disease segmentation. In the classification task, we categorize approaches into basic deep learning models, transfer learning, fine-tuning, and custom models, emphasizing how these advanced techniques have progressively improved diagnostic outcomes.
@@ 2.1.2 U-Net Enhancements with Transformers
Recent research has significantly advanced lung segmentation by enhancing the U-Net architecture with attention mechanisms. Oktay et al. oktay2018attention introduced mechanisms that enable the model to concentrate on the most crucial areas within chest X-rays using Attention Gates (AGs). This innovation enhances segmentation accuracy and sensitivity to disease characteristics. Additionally, research by Wu et al. wu2023transformer , Gu et al. gu2018recent , and Liu et al. liu2024sta has shown that incorporating attention mechanisms into the U-Net framework significantly improves lung segmentation performance, underscoring the effectiveness of this method in precisely outlining lung boundaries and enhancing diagnostic outcomes.
@@ 2.2.2 Deep Learning Models
The beginning of deep learning, especially Convolutional Neural Networks (CNNs), has significantly transformed medical image analysis by providing superior accuracy and robustness. For instance, Stephen et al. designed a custom CNN model from scratch, achieving a training accuracy of 95.31% and a validation accuracy of 93.73% stephen2019efficient . Similarly, Sharma et al. created a straightforward CNN architecture that reached a 90.68% accuracy rate on the ”Kermany” dataset using data augmentation sharma2020feature ; kermany2018labeled . However, relying solely on data augmentation does not introduce substantially new information, restricting the model’s ability to learn a wide range of complex patterns from the training data.
@@ 3.1 Overview
Our approach begins with a segmentation step where we employ a TransUNet model, which integrates transformer-based attention mechanisms into the popular U-Net architecture. This model is trained on ”Chest Xray Masks and Labels” dataset candemir2013lung ; jaeger2013automatic to accurately segment lung regions in the images. By predicting masks for ”Cohen” dataset cohen2020covid using this pre-trained TransUNet, we can isolate the regions of interest, enhancing the subsequent classification task.
The segmentation step provides us with precise lung masks, ensuring that our classification model focuses on the relevant areas of the X-ray images. This preprocessing step is crucial for improving the overall accuracy of the system by reducing background noise and irrelevant features.
Our classification approach extracts multi-scale feature maps from three key stages of the ResNet models: the outputs of Block 2, Block 3, and Block 4. These stages provide a rich set of features at different scales, which are crucial for accurately identifying Pneumonia in chest X-rays. The extracted feature maps are then processed through a specialized transformer module that employs a cross-attention mechanism. This transformer enhances the feature representation by allowing the network to focus on the most relevant parts of the image.
After the cross-attention processing, the feature maps are concatenated to form a comprehensive representation of the input image. This combined feature map is subsequently fed into fully connected layers to perform the final classification. The overall architecture is designed to effectively integrate multi-scale information and attention mechanisms, thereby improving the classification accuracy. An overview of the proposed method is illustrated in Figure 1.
@@ 3.2.1 TransUNet Architecture
The decoder aims to reconstruct the segmented output by progressively upsampling the feature maps and concatenating them with corresponding encoder layers. Each upsampling step is followed by a double convolution block, which refines the features and reduces the number of channels. This structure helps in restoring the spatial resolution of the feature maps while retaining the detailed information captured by the encoder. The final layer of the decoder is a convolutional layer with a single output channel, which predicts the segmentation mask for the lung regions.
@@ 3.3.1 Backbone
The backbone of our proposed method utilizes a pre-trained ResNet model, specifically ResNet-50 or ResNet-101, to extract multi-scale feature maps from the input chest X-ray images. Initially, we resize all images to 512x512 pixels. We focus on the outputs from Blocks 2, 3, and 4 of the ResNet, denoted as $B^{2}$, $B^{3}$, and $B^{4}$ respectively. Each of these blocks provides feature maps of size $(c,h,w)$, where $h=w=64$ The channels $c$ for these blocks are 512, 1024, and 2048 respectively.

### v3

@@ Abstract
Pneumonia, a prevalent respiratory infection, remains a leading cause of morbidity and mortality worldwide, particularly among vulnerable populations. Chest X-rays serve as a primary tool for pneumonia detection; however, variations in imaging conditions and subtle visual indicators complicate consistent interpretation. Automated tools can enhance traditional methods by improving diagnostic reliability and supporting clinical decision-making. In this study, we propose a novel multi-scale transformer approach for pneumonia detection that integrates lung segmentation and classification into a unified framework. Our method introduces a lightweight transformer-enhanced TransUNet for precise lung segmentation, achieving a Dice score of 95.68% on the ”Chest X-ray Masks and Labels” dataset with fewer parameters than traditional transformers. For classification, we employ pre-trained ResNet models (ResNet-50 and ResNet-101) to extract multi-scale feature maps, which are then processed through a modified transformer module to enhance pneumonia detection. This integration of multi-scale feature extraction and lightweight transformer modules ensures robust performance, making our method suitable for resource-constrained clinical environments. Our approach achieves 93.75% accuracy on the ”Kermany” dataset and 96.04% accuracy on the ”Cohen” dataset, outperforming existing methods while maintaining computational efficiency. This work demonstrates the potential of multi-scale transformer architectures to improve pneumonia diagnosis, offering a scalable and accurate solution to global healthcare challenges. https://github.com/amirrezafateh/Multi-Scale-Transformer-Pneumonia
@@ 1 Introduction
Pneumonia is a serious respiratory condition that causes inflammation in one or both lungs, leading to symptoms such as fever, cough, and difficulty breathing. This illness is particularly dangerous for young children, accounting for approximately 15% of mortality in children under the age of five unicef_pneumonia . The disease is more prevalent in developing countries, where limited access to healthcare, pollution, overcrowding, and poor living conditions exacerbate its effects who_pneumonia_in_children .
Early and accurate diagnosis is essential for effective treatment; however, pneumonia can be challenging to identify due to its similarity to other lung diseases health_poverty_action_poverty_and_health . Chest X-rays are commonly used for diagnosis due to their cost-effectiveness and non-invasive nature radiologyinfo_chest_xray . Nevertheless, the interpretation of these images can vary significantly, underscoring the necessity for consistent and automated diagnostic tools.
Recent advancements in deep learning, particularly in Convolutional Neural Networks (CNNs), have shown significant promise in improving pneumonia diagnosis from chest X-rays yan2023pneumonia ; lin2023enhancing . These models can analyze medical images with remarkable precision, often outperforming human radiologists in both consistency and speed. Recent innovations, such as attention mechanisms, have further enhanced diagnostic accuracy askari2024enhancing ; parhami2023comparison ; fateh2024advancing . Additionally, transformers have demonstrated significant potential in medical imaging tasks due to their ability to model long-range dependencies and identify complex patterns aurangzeb2024introducing ; fateh2024msdnet . These advancements highlight the potential of AI to complement radiologists’ expertise and enhance patient outcomes.
Lung segmentation is a critical preprocessing step for improving the accuracy of pneumonia detection in chest X-rays. However, this task faces several challenges. First, the presence of artifacts, overlapping anatomical structures, and low contrast in chest X-rays can make it difficult to accurately delineate lung boundaries alam2024attention . Traditional segmentation models like U-Net, while effective, often struggle to capture fine-grained details and contextual information, leading to suboptimal performance in complex cases zhang2023cams . Additionally, the high variability in lung shapes and sizes across patients further complicates the segmentation process. Although transformer-based models have shown promise in addressing these issues, they often come with high computational costs and large parameter counts, making them unsuitable for resource-constrained environments cao2022swin .
Classification of pneumonia from chest X-rays also presents significant challenges. First, the subtle visual indicators of pneumonia, such as small opacities or localized consolidations, can be easily missed by traditional CNN-based models like ResNet and DenseNet, which primarily focus on local features rajpurkar2017chexnet . While transformer-based models excel at capturing global contextual information, they require large amounts of labeled data and extensive computational resources, limiting their practicality in real-world clinical settings rezvani2024fusionlungnet . Moreover, the domain shift between pre-trained models (e.g., on ImageNet) and medical imaging datasets often results in suboptimal performance, necessitating advanced techniques like transfer learning and domain adaptation chen2017rethinking . These challenges highlight the need for a computationally efficient and accurate model that can effectively leverage both local and global features for pneumonia diagnosis.
@@
To address these challenges, we propose an innovative approach that leverages deep learning with an integrated lightweight transformer, significantly reducing the number of parameters compared to traditional transformers while maintaining lower model complexity. Our method begins with lung segmentation using a TransUNet model, which integrates transformer-based attention mechanisms into the U-Net architecture. The TransUNet model is trained on the ”Chest Xray Masks and Labels” dataset candemir2013lung ; jaeger2013automatic to accurately segment lung regions in the images. Once trained, this pre-trained model is used with frozen weights to predict lung masks for our target datasets, ”Kermany” kermany2018labeled and ”Cohen” cohen2020covid . This segmentation step isolates the lung regions, enhancing the subsequent classification task.
@@
  • Introduction of a novel TransUNet architecture for the segmentation task, achieving a Dice score of 95.68% on the ”Chest X-ray Masks and Labels” dataset.
@@
  • Achieving high accuracy rates of 93.75% on the ”Kermany” dataset and 96.04% on the ”Cohen” dataset.
@@ 2 Related Work
In recent years, the focus of research on diagnosing and categorizing lung diseases, including pneumonia, through medical imaging has intensified, driven by advances in machine learning and deep learning technologies jennifer2023neutrosophic . Precisely segmenting lung areas in chest X-ray (CXR) images is essential for reliable disease identification and thorough analysis. This section examines deep learning techniques for segmenting and diagnosing lung diseases in chest X-ray (CXR) images. For the segmentation task, we focus on the U-Net architecture and its variations, including attention mechanisms and transformer blocks, which have significantly advanced lung disease segmentation. In the classification task, we categorize approaches into basic deep learning models, transfer learning, fine-tuning, and custom models, emphasizing how these advanced techniques have progressively improved diagnostic outcomes.
@@ 2.1.1 U-Net for CXR Segmentation
However, traditional U-Net architectures face several limitations that impede their effectiveness in complex segmentation tasks. These limitations include the inability to leverage multi-scale information, which is essential for capturing fine-grained details, and difficulties in extracting rich contextual information, particularly for small or complex anatomical structures huang2020unet . Furthermore, the simple skip connections in U-Net may transfer irrelevant or noisy features, leading to ambiguity in feature representation and reduced segmentation accuracy jha2020doubleu . These challenges are especially problematic in chest X-rays, where overlapping structures and low contrast exacerbate noise.
@@ 2.1.2 U-Net Enhancements with Transformers
To address traditional U-Net limitations, advanced architectures that enhance U-Net’s ability to capture multi-scale and contextual information are needed. Recent research has significantly advanced lung segmentation by enhancing the U-Net architecture with attention mechanisms. Oktay et al. oktay2018attention introduced mechanisms that enable the model to concentrate on the most crucial areas within chest X-rays using Attention Gates (AGs). This innovation enhances segmentation accuracy and sensitivity to disease characteristics. Additionally, research by Wu et al. wu2023transformer , Gu et al. gu2018recent , and Liu et al. liu2024sta has shown that incorporating attention mechanisms into the U-Net framework significantly improves lung segmentation performance, underscoring the effectiveness of this method in precisely outlining lung boundaries and enhancing diagnostic outcomes.
@@ 2.2.2 Deep Learning Models
The beginning of deep learning, especially Convolutional Neural Networks (CNNs), has significantly transformed medical image analysis by providing superior accuracy and robustness sanghvi2023deep ; shiri2024differentiation rajinikanth2022pneumonia ; allioui2022multi . For instance, Stephen et al. designed a custom CNN model from scratch, achieving a training accuracy of 95.31% and a validation accuracy of 93.73% stephen2019efficient . Similarly, Sharma et al. created a straightforward CNN architecture that reached a 90.68% accuracy rate on the ”Kermany” dataset using data augmentation sharma2020feature ; kermany2018labeled . Rehman et al rehman2021self , used self-activated neural networks for multi-class classification to detect chest diagnosis. However, relying solely on data augmentation does not introduce substantially new information, restricting the model’s ability to learn a wide range of complex patterns from the training data.
@@ 2.2.4 Ensemble Approaches
Despite their success, pre-trained models such as ResNet have inherent limitations. While ResNet models are powerful, they often struggle to independently capture all the discriminative features required for specific tasks, particularly in complex medical imaging scenarios like pneumonia detection jaiswal2019identifying . This limitation is evident in studies where ResNet architectures require complementary support from other models or advanced techniques, such as snapshot ensembling and weighted averaging, to achieve optimal performance gabruseva2020deep . Furthermore, ResNet’s reliance on local feature extraction through convolutional layers can hinder its ability to model long-range dependencies and global contextual information, which are crucial for accurate classification in medical images chen2021transunet . These shortcomings highlight the need for more robust frameworks, such as transformers, which excel at capturing global context and intricate patterns, thereby addressing the limitations of traditional CNN-based models like ResNet.
@@ 2.2.5 Transformers
In recent years, transformer-based models have continued to evolve, with a particular emphasis on improving efficiency and accuracy in medical imaging tasks. The Swin Transformer V2 liu2022swin has emerged as a powerful architecture for various medical imaging tasks, including pneumonia detection. It achieves superior performance by leveraging hierarchical feature extraction and shifted window mechanisms, which allow it to capture both local and global patterns in chest X-rays. In a recent study, the Swin Transformer V2 achieved an accuracy of 98.6% on a diverse chest X-ray dataset, outperforming traditional CNNs like ResNet and DenseNet mishra2024empowering . This highlights its potential for clinical applications where high diagnostic accuracy is essential.
Hybrid architectures that combine CNNs and transformers have demonstrated remarkable success in pneumonia detection. For instance, a hybrid model integrating ResNet34 with a Multi-Axis Vision Transformer achieved a state-of-the-art accuracy of 94.87% on the Kaggle pediatric pneumonia dataset. This model leverages the local feature extraction capabilities of CNNs and the global context modeling of transformers, resulting in fewer misclassifications and improved robustness angara2024novel . Additionally, another study proposed a hybrid CNN-Swin Transformer model that achieved 98.72% accuracy, demonstrating the effectiveness of combining convolutional layers with transformer blocks for enhanced feature extraction mustapha2025enhanced .
While transformer-based models have shown significant promise in medical image classification, they face several limitations when applied to pneumonia detection. First, many existing transformer architectures, such as Vision Transformers (ViTs) and Swin Transformers, require large computational resources and extensive training data, making them unsuitable for resource-constrained clinical environments liu2022swin . Additionally, these models often struggle to effectively combine local feature extraction (a strength of CNNs) with global context modeling (a strength of transformers), leading to suboptimal performance in tasks like pneumonia detection, where both fine-grained details and global patterns are critical angara2024novel . Furthermore, the high parameter counts and complexity of traditional transformers can result in longer training times and higher hardware requirements, limiting their practicality in real-world applications mustapha2025enhanced .
@@ 3.1 Overview
Our approach begins with a segmentation step where we employ a TransUNet model, which integrates transformer-based attention mechanisms into the popular U-Net architecture. This model is trained on ”Chest Xray Masks and Labels” dataset candemir2013lung ; jaeger2013automatic to accurately segment lung regions in the images. By predicting masks for ”Cohen” dataset cohen2020covid using this pre-trained TransUNet, we can isolate the regions of interest, enhancing the subsequent classification task. The segmentation step provides us with precise lung masks, ensuring that our classification model focuses on the relevant areas of the X-ray images. This preprocessing step is crucial for improving the overall accuracy of the system by reducing background noise and irrelevant features.
Our classification approach extracts multi-scale feature maps from three key stages of the ResNet models: the outputs of Block 2, Block 3, and Block 4. These stages provide a rich set of features at different scales, which are crucial for accurately identifying Pneumonia in chest X-rays. The extracted feature maps are then processed through a specialized transformer module that employs an attention mechanism. This transformer enhances the feature representation by allowing the network to focus on the most relevant parts of the image.
After the attention processing, the feature maps are concatenated to form a comprehensive representation of the input image. This combined feature map is subsequently fed into fully connected layers to perform the final classification. The overall architecture is designed to effectively integrate multi-scale information and attention mechanisms, thereby improving the classification accuracy. An overview of the proposed method is illustrated in Figure 1.
@@ 3.2.1 TransUNet Architecture
Also, the transformer modules integrate into each skip connection between the encoder and decoder. These transformers enhance the model’s ability to capture global contextual information at each resolution level. In this design, the query for each transformer’s attention mechanism is derived from the output of the transformer at the preceding, lower level. The transformer’s output is then element-wise multiplied with the original skip connection feature maps before being passed to the decoder. This ensures that the information passed to the decoder not only retains local details but also incorporates refined global representations.
The decoder reconstructs the segmented output by progressively upsampling the feature maps and concatenating them with the outputs of the transformer-augmented skip connections. Each upsampling step is followed by a double convolution block to refine the features and reduce the number of channels. This structure allows the decoder to restore the spatial resolution of the feature maps while retaining the detailed information captured by the encoder. Finally, a convolutional layer with a single output channel generates the segmentation mask for the lung regions.
@@ 3.3.1 Backbone
The backbone of our proposed method utilizes a pre-trained ResNet model, specifically ResNet-50 or ResNet-101, to extract multi-scale feature maps from the input chest X-ray images. Initially, we resize all images to 512x512 pixels. We focus on the outputs from Block 2, Block 3, and Block 4 of the ResNet, denoted as $B^{2}$, $B^{3}$, and $B^{4}$ respectively. Each of these blocks provides feature maps of size $(c,h,w)$, where $h=w=64$ The channels $c$ for these blocks are 512, 1024, and 2048 respectively.
The selection of Block 2, Block 3, and Block 4 is motivated by their unique contributions to the classification task. Block 2 captures low-level features such as edges and textures, which are essential for identifying subtle patterns in the lung regions. Block 3 extracts mid-level features, including more complex structures like lung lobes and localized consolidations, which are critical for detecting early signs of pneumonia. Block 4 provides high-level semantic features, such as global contextual information and disease-specific patterns, which are necessary for accurate classification. By combining these multi-scale features, our model can effectively capture both fine-grained details and global context, leading to improved diagnostic performance.
@@ 4.2 Experimental Setting
4.3 Evaluation Metrics
Accuracy=TP+TNTP + TN + FP + $\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP + TN + FP + FN}}$ (10)
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{TN}}$ (11)
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (12)
[…]
MCC: The Matthews Correlation Coefficient (MCC) is a robust metric for evaluating segmentation performance, especially in imbalanced datasets. It considers true positives (TP), true negatives (TN), false positives (FP), and false negatives (FN) to provide a balanced measure of classification quality. As stated in anbalagan2023analysis , MCC is particularly useful for tasks where class imbalance is a concern, as it accounts for all four categories of the confusion matrix. MCC is calculated using Equation 14.
$\text{MCC}=\frac{\text{TP}\times\text{TN}-\text{FP}\times\text{FN}}{\sqrt{(\text{TP}+\text{FP})\times(\text{TP}+\text{FN})\times(\text{TN}+\text{FP})\times(\text{TN}+\text{FN})}}$
Dice Coefficient: The Dice Coefficient, also known as the F1-score for segmentation tasks, measures the overlap between the predicted segmentation and the ground truth. It is particularly useful for evaluating the accuracy of region-based segmentation, such as lung segmentation in chest X-rays. The Dice Coefficient is calculated using Equation 15:
$\text{Dice}=\frac{2\times\text{TP}}{2\times\text{TP}+\text{FP}+\text{FN}}$ (15)
4.4 Comparison with State-of-the-Art
Table 1: Performance on ”Chest X-ray Masks and Labels” dataset for segmentation task. Numbers in bold represent the best performance, while underlined values denote the second-best performance.
Model Dice Accuracy Precision Recall F1-score MCC
Unet ronneberger2015u 93.46% 96.88% 97.41% 89.83% 93.46% 91.55%
RU-Net alom2018nuclei 92.07% 96.34% 99.58% 85.61% 92.07% 90.13%
ResNet34-Unet lau2020automated 93.83% 97.06% 98.13% 89.89% 93.83% 92.06%
BCDU-Net azad2019bi 94.14% 97.20% 98.25% 90.37% 94.14% 92.44%
ResBCDUnet jalali2021resbcdu 94.34% 97.31% 98.89% 90.20% 94.34% 92.75%
NasNet zhang2022automatic 94.95% 97.52% 96.55% 93.42% 94.95% 93.33%
DABT-U-Net jalali2024dabt 95.11% 97.64% 98.25% 92.16% 95.11% 93.64%
ABANet rezvani2024abanet 95.25% 97.71% 98.53% 92.18% 95.25% 93.84%
FusionLungNet rezvani2024fusionlungnet 95.29% 97.73% 98.66% 92.14% 95.29% 93.89%
Our 95.7%$\pm$0.5 97.9%$\pm$0.5 97.5%$\pm$0.6 93.9%$\pm$0.6 95.7%$\pm$0.5 94.3%$\pm$0.6